145.3 g of Sodium phosphate reacted completely with Calcium chloride. The reaction yielded to new compounds.

(note: Na=23.0g/mol, P=31.0g/mol, O=16.0g/mol, Ca=40.0g/mol, Cl=35.0g/mol)

Please find the compounds.


1
Expert's answer
2020-12-30T04:02:54-0500

3CaCl2+ 2Na3PO4= Ca3(PO4)+ 6NaCl

calcium phosphate, sodium chloride

find the masses of the obtained compounds:

"n (Na3PO4)=\frac{m}{M}=\frac{145.3}{164}=0.89"

masses ofCa3(PO4)2

"\frac{0.89}{2}=\frac{x}{1}"

x=0.445

"m=M\times n=0.445*310=137.95"

masses of NaCl:

"\frac{0.89}{2}=\frac{x}{6}"

x=2.67

"m=M\times n=2.67*58=154.86"